Raymond Mining Corporation has 10.0 million shares of common stock outstanding, 440,000 shares of 4% $100 par value preferred stock outstanding, and 173,000 7.50% semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $48 per share and has a beta of 1.50, the preferred stock currently sells for $97 per share, and the bonds have 10 years to maturity and sell for 115% of par. The market risk premium is 8.8%, T-bills are yielding 5%, and Raymond Mining’s tax is 40%.

a. What is the firm’s market value capital structure?

Market value
Debt
Equity
Preferred stock

b. If Raymond Mining is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 3 decimal places.)
Answer #1

1.
=173000*1000*115%/(173000*1000*115%+10*10^6*48+440000*97)=0.275695300915982

2.
=10*10^6*48/(173000*1000*115%+10*10^6*48+440000*97)=0.665160816484902

3.
=440000*97/(173000*1000*115%+10*10^6*48+440000*97)=0.0591438825991159

4.
=(173000*1000*115%*(RATE(10*2,7.5%*1000/2,-115%*1000,1000)*2*(1-40%))+10*10^6*48*(5%+1.5*8.8%)+440000*97*(4%*100/97))/(173000*1000*115%+10*10^6*48+440000*97)=13.2641346915888%